The length of a rectangle is 6 cm more than 1 1/2 times the width. The perimiter of the rectangle is 56 cm. Find the dimensions of the rectangle.Can someone show me the work too?
Answer by rfer(16322)

width=x=8.8
length=1.5x+6=19.2
---------------
P=2L+2W
2(1.5x+6)+2x=56
3x+12+2x=56
subtract 12 from both sides.
5x=56-12
5x=44
x=8.8
--------------
P=2(8.8)+2(19.2)=56